# HPC-Driven Modeling with ML-Based Surrogates for Magnon-Photon Dynamics in Hybrid Quantum Systems

### v1

@@ Abstract
Simulating hybrid magnonic quantum systems remains a challenge due to the large disparity between the timescales of the two systems. We present a massively parallel GPU-based simulation framework that enables fully coupled, large-scale modeling of on-chip magnon-photon circuits. Our approach resolves the dynamic interaction between ferromagnetic and electromagnetic fields with high spatiotemporal fidelity. To accelerate design workflows, we develop a physics-informed machine learning surrogate trained on the simulation data, reducing computational cost while maintaining accuracy. This combined approach reveals real-time energy exchange dynamics and reproduces key phenomena such as anti-crossing behavior and the suppression of ferromagnetic resonance under strong electromagnetic fields. By addressing the multiscale and multiphysics challenges in magnon-photon modeling, our framework enables scalable simulation and rapid prototyping of next-generation quantum and spintronic devices.
@@ 1 Introduction
Hybrid quantum systems, which combine distinct physical platforms, are a promising route toward advanced quantum technologies, as they harness strong interactions that may not be readily achievable in a single platform [1, 2]. These systems take many forms, coupling any two (or more) quantum platforms — for example, superconducting qubits  [3, 4], microwave resonators  [5], single spins [6], spin ensembles  [4, 7, 8, 9], or mechanical resonators [10, 11, 12] — to harness strong interactions. These heterogeneous systems leverage complementary advantages of each component, but their rich multi-physics interactions pose formidable modeling challenges. A prominent example is cavity magnonics, where collective spin excitations (magnons) couple with microwave photons in a resonant cavity to form hybrid magnon-polariton modes when tuned into resonance [13, 14, 15]. While an avoided crossing indicates hybridization, it is not sufficient for strong coupling; the system enters the strong-coupling regime only when the coupling rate $g$ (half the on‑resonance splitting) exceeds the dissipation rates of the uncoupled magnon and photon modes, in which case energy oscillates coherently between them and the hybrid modes exhibit remarkable coherence owing to the high $Q$ factors of microwave resonators and the narrow ferromagnetic resonance (FMR) linewidth [16]. These states are essential for quantum operations such as mode swapping [16, 17], quantum state storage  [18, 4, 19], and dynamic control of energy exchange  [20, 19]. The hallmark experimental signature of strong magnon–photon coupling is a pronounced avoided crossing (mode splitting) in the frequency spectrum, in agreement with theoretical predictions [21] and observed in many 3D  [13, 22] and on-chip 2D [23, 7, 8] cavity based systems. In the time domain, the coupled system exhibits Rabi-like oscillations with very high energy exchange per cycle, reflecting the complete hybridization of the two excitations into a single quantum system [17]. Achieving practical quantum operations with magnon-photon systems requires precise dynamical control over energy transfer between the two hybrid magnon-polariton modes [24, 25]. Therefore, developing quantum devices that leverage these effects necessitates a detailed understanding of their time-domain dynamics.
@@
However, accurately modeling the complex dynamics of such hybrid systems presents a significant computational challenge. Existing modeling approaches have been limited in scope. Most analytical treatments require inputting an assumed magnon-photon coupling strength, and the few numerical studies to date have often resorted to reduced dimensions (1D or 2D models) or macrospin approximations to mitigate computational cost [21]. Conventional micromagnetic modeling strategies rely on the magnetostatic approximation [26, 27, 28], which has provided valuable insights, but which typically neglects the dynamic nature of electromagnetic (EM) fields, including time-dependent displacement currents. The underlying physics of magnon-photon coupling spans multiple domains and scales, requiring a multiphysics simulation that solves Maxwell’s equations for electromagnetic fields alongside the Landau-Lifshitz-Gilbert (LLG) equation for magnetization dynamics [17]. Solving these coupled equations self-consistently in three dimensions with realistic material parameters is computationally intensive, often necessitating high-performance computing (HPC) resources to achieve adequate spatial and temporal resolution [29]. The disparate time scales and length scales involved – from microwave photons to nanometer-scale spin precessions – further complicate direct simulation. Maxwell-LLG coupling is achieved by self-consistently updating the dynamic EM fields in sync with magnetization dynamics, thereby capturing the bidirectional interaction between microwave photons and magnons  [29, 30, 31, 17, 32, 33, 34, 35]. The majority of such time-domain dynamical simulations employ either explicit schemes like finite-difference time-domain (FDTD) leapfrog integrators  [29, 34, 17, 33, 32], or implicit schemes such as alternating-direction implicit (ADI) FDTD and Crank-Nicolson methods  [30, 31]. While explicit FDTD is straightforward, it is only conditionally stable (restricted by the Courant limit), leading to significant oversampling of the EM wave in the time domain. To ease numerical constraints, researchers have explored strategies such as artificially increasing the medium permittivity – an approach that relaxes time – step oversampling [17]. However, this strategy might be difficult to extend to complex device structures with highly inhomogeneous local permittivity and conductivities, as in on-chip devices. On the other hand, implicit methods offer improved (often unconditional) stability, but they require solving more complex update steps. Additionally, implementing precise boundary conditions in implicit mathematical frameworks, such as perfectly matched layers (PML), is challenging and limits scalability in real-world quantum devices  [31]. HPC techniques, including GPU acceleration and parallel FDTD-LLG solvers combined with domain decomposition, significantly enhance simulation speed and scalability, enabling large-scale modeling on modern supercomputing architectures  [36, 29]. Finite-element time-domain models that rely on commercial software have also been adopted, but they suffer from the overhead of assembling and solving large linear systems at each time step, leading to high CPU and memory usage [35]. Nonetheless, to the best of the authors’ knowledge, dynamical time-domain simulation of practical sized, on-chip hybrid magnon-photon circuits has remained elusive [7, 8]. This underscores the need for more powerful computational methods to simulate hybrid magnon–photon systems in full detail.
@@
The main numerical challenges arise from the need for large-scale three-dimensional (3D) simulations and fine meshes to accurately capture the detailed structures. In many practical scenarios, however, only a few spatial observation points are needed for circuit analysis tasks. This discrepancy underscores the potential for more efficient modeling strategies that focus resources on critical regions and time windows, reducing overall costs while preserving key physics. Machine learning (ML) has emerged as a powerful approach to tackle the computational challenges posed by high-dimensional and nonlinear dynamical systems [37, 38]. In particular, frameworks such as Physics-Informed Neural Networks (PINNs) [39, 40], Neural Operators (e.g., DeepONet  [41] and Fourier Neural Operator [42]), and Neural Discrete Equilibrium (NeurDE) [43] have demonstrated strong capabilities in modeling partial differential equations (PDEs), by efficiently capturing their underlying spatial-temporal dynamics. Furthermore, recent advances in sequence modeling architectures, including Recurrent Neural Networks (RNNs) [44], state-space models (SSMs) [45, 46] and Transformers [47], have shown remarkable success in predicting complex time-evolving physical phenomena [48, 49, 50, 51, 52, 53, 54, 55].
@@ 2 Results
Figure 1: Magnon-photon hybrid quantum circuit modeling setup and simulated field profiles. (a) Perspective view of the magnon-photon coupling circuit, comprising a CPW resonator and a ferromagnetic thin film. The CPW resonator consists of a center conductor and lateral ground conductors (yellow) positioned on a silicon substrate (blue). The ferromagnet, shown as a red stripe, is placed on top of the center conductor. Material and geometrical parameters are detailed in the main text. (b)-(d) Cross-sectional views of the circuit, overlaid with the simulated magnetic field (H) distribution at different locations along the longitudinal direction of the CPW resonator, as indicated in (a). (e)-(g) Top views of the circuit, overlaid with the simulated electric and magnetic field distributions in the z-plane at different observation points, as marked in (a). (h) Theoretical field distributions along the longitudinal direction of the circuit, where the electric field $E_{x}$ is concentrated in the lateral air gap, and the magnetic field $H_{x}$ is observed above and beneath the center conductor.
@@
We employ a fully coupled dynamic Maxwell-LLG solver, ARTEMIS, based on the approach in [29], using an explicit FDTD leap-frog scheme to update the electric field E, magnetic field H, and magnetization M. In nonmagnetic regions, only Maxwell’s equations are solved (with M  set to zero). In magnetic regions, Maxwell’s equations and the LLG equation are solved self-consistently in a mutual feedback loop (see Numerical Model. in Methods for details). All field variables computed by the solver – E, H, and M – achieve second-order accuracy, owing to the trapezoidal time-domain discretization of the LLG equation and the leap-frog time-marching scheme for Maxwell’s equations. To handle large problem sizes, we implement both multicore and GPU parallelization for the spatial mesh, as illustrated in Figure 2(a). We use the AMReX [56] library to partition the computational domain into non-overlapping grids, each assigned to an MPI rank. AMReX follows an MPI+X parallelization model, where X can be OpenMP (for multicore CPUs) or CUDA (for GPUs).
@@
The input data for the physics-informed ML model is collected from nine discrete spatial locations, shown in Figure 2(b). These points lie in the middle of the ferromagnet layer along its thickness and include the film’s center, as well as the four corners and midpoints of the four edges in the lateral plane. More details can be found in the ML Training Implementation. section of Methods. Figure 2(f) illustrates the general ML framework for sequential modeling of the magnetization. The model is based on the Long Expressive Memory (LEM) module [57], and it comprises an LEM encoder and decoder. The structure of the LEM cell is illustrated in Figure 2(d), which shows the computation flow across its internal states. At each time step, the input $X_{t}$, the previous hidden state $H_{t-1}$ and latent state $C_{t-1}$ are linearly projected and combined to generate two adaptive time-step parameters: $\Delta t$ for updating the latent state $C_{t}$, and $\overline{\Delta t}$ for updating the hidden state $H_{t}$. These two time-step parameters act as gating coefficients, controlling the extent to which newly computed nonlinear activations (via $\tanh$) are blended into each state. The latent state $C_{t}$ is updated first, capturing fast-evolving features from the input and hidden state. This updated $C_{t}$ is then transformed and used to update the hidden state $H_{t}$, which serves as the main memory for downstream prediction. Additional implementation details and the mathematical formulation are provided in the LEM Cell. of Methods. The LEM module processes the input sequence (200 temporal samples) using a stacked encoder composed of two LEM cells, which evolve coupled hidden $H$ and latent states $C$ with adaptive time constants. The final states from the encoder are used to initialize a two-stage decoder, which begins from a zero input and autoregressively generates a fixed-length output sequence. At each decoding step, the internal states are updated using the same LEM dynamics, and the top-layer hidden state is passed through a fully connected layer to produce the output. The model is trained using a composite loss function that integrates three component:
@@
Given the complexity of the hybrid quantum system, we incorporated Curriculum Learning [58] to improve generalization and help the model better capture long-term dynamics and the underlying physics. As shown in Figure 2(e), this approach enables the model to handle tasks of increasing complexity over time. During training, three key hyperparameters are dynamically adjusted:
@@
Training begins with simpler scenarios: shorter prediction sequences, no physics loss, and a higher learning rate, and gradually moves to more complex cases involving longer sequences, physics loss, and a lower learning rate. This staged approach ensures that the model generalizes effectively while capturing the intricate physics governing the field interactions. Details on how the training and testing datasets were curated are provided in the Datasets Curation. section of Methods.
Figure 2: The HPC and ML frameworks for magnon-photon hybrid quantum modeling. (a) Simulation domain setup and GPU partitioning in the Maxwell-LLG coupled numerical model, ARTEMIS. The GPU parallelization leverages the grid-structured GPU library AMReX [56], achieving nearly flat weak-scaling performance. (b) Top view of the magnon-photon hybrid quantum circuit; the cross marks indicate probing points used for data collection. (c) Example time-series data of the magnetization M recorded at a single probing point. (d) Detailed architecture of the LEM cell. The cell updates a latent state ($C_{t}$) and a hidden state ($H_{t}$) using the input $X_{t}$, previous states, and learnable time constants ($\Delta t$). $C_{t}$ captures short-term dynamics, while $H_{t}$ integrates transformed latent features to maintain temporal context. (e) The curriculum learning strategy used to train the ML model, with a staged schedule with progressively adjusted learning rates, batch sizes, sequence lengths, and physics-based loss weights to stabilize training and enhance convergence. (f) Overall ML framework for time-series prediction, built out of the LEM-based architecture. The encoder consists of sequential LEM Cells that process the input sequence (X) into two evolving internal states: the latent state (C); and the hidden state (H). These states are used to initialize an auto-regressive decoder, which also consists of stacked LEM Cells. At each step, the decoder outputs a hidden state (H) that is passed through a shared fully connected (FC) layer to produce the predicted output.
@@
Our Maxwell-LLG dynamical modeling has predicted the dynamically evolving and interacting electric field and magnetization, shown in Figure 3. Figure 3(a) presents top-view snapshots of $E_{x}$ (top row) and $M_{x}$ (bottom row) at selected time instants, with an applied magnetic DC bias $H_{0}=2050$ Oersted. According to the Kittel’s law, under such $H_{0}$, the FMR frequency is $f_{\textrm{m}}=\omega_{\textrm{m}}/2\pi=(\gamma/2\pi)\sqrt{H_{0}(H_{0}+4\pi M_{S})}=15.027$ GHz, detuned from the photon frequency $f_{\textrm{p}}$ by about $\Delta=200$ MHz. This small detune enables strong mode hybridization and, consequently, strong magnon-photon coupling. The $E_{x}$ fields are observed at the same location described in Figure 1(e), i.e., along the vertical center plane passing through the CPW resonator’s lateral air gap. The $M_{x}$ fields are observed at the same location described in Figure 1(f), i.e., on the top surface of the ferromagnetic stripe. For additional details on these observation points, see Figure 1(a). At 0.0235 ns, the electric field appears only in the feeding patches, without exciting resonant modes (photons) in the CPW resonator, and no dynamic magnetization is observed in the ferromagnet. By 0.3528 ns, both $E_{x}$ and $M_{x}$ emerge in the resonator. Starting at 0.9819 ns, even though the input field excitation has vanished (the input signal completely decays by 0.5 ns, as shown in Figure 3(b)), oscillations of $E_{x}$ and $M_{x}$ persist within the CPW resonator. Notably, $E_{x}$ peaks near the resonator’s ends, whereas $M_{x}$ peaks at the center. Both fields exhibit well-defined oscillations with only minor decay, indicating low damping, which is a result of minimal Gilbert damping, small metallic losses, and limited electromagnetic leakage into air or substrate. Furthermore, $E_{x}$ and $M_{x}$ alternate in reaching their standing-wave maxima: when $E_{x}$ is at a peak, $M_{x}$ is at a node, and vice versa. This phase alternation underscores strong energy exchange between the magnon and photon polaritons, more clearly seen in the transient response of Figure 3(c), and the yielded mode splitting in the Fourier transformed $E_{x}$ spectra shown in Figure 3(d). Figure 3(e) visualizes the spectra of $E_{x}$ magnitude under various $H_{0}$, revealing a clear anti-crossing behavior, indicating coherent coupling between magnons and photons [14]. The simulated spectrum matches theoretical analysis as shown in Extended Data Figure 1. The $S$-parameter ($|S_{12}|$) spectra shown in Figure 3(f) is converted from the $E_{x}$ field following the workflow based on field calculation in  [59], indicating the transmission rate from one end to the other of the microwave resonator, also showing an anti-crossing spectra. Both figures show mode splitting $g_{\textrm{mp}}=670$ MHz, where $g_{\textrm{mp}}$ is the total magnon-photon coupling strength, defined as $g_{\textrm{mp}}=g_{0}\sqrt{N}$, with $g_{0}$ defined as the coupling strength between photons and individual spins in the ferromagnet, and $N$ is the number of spins in the ferromagnet [21, 13]. $g_{0}$ contributes to the total Hamiltonian of the system:
@@
One unique advantage of numerical simulation is the ability to generate signals on demand, including those with intensities that are experimentally unattainable. Extended Data Figure 2 displays the electric-field spectra as the excitation intensity is increased. It demonstrates that, in the absence of significant spin-spin interactions (e.g., exchange coupling), progressively stronger microwave magnetic fields suppress the magnon dynamics and leave only the cavity photon mode. Such behaviour arises because, at high powers, nonlinear spin-wave interactions (particularly three-magnon splitting) induce a Suhl instability that transfers energy from the uniform ferromagnetic resonance to non-uniform spin waves; this saturation of the Kittel mode causes the magnon-photon anticrossing gap to close and yields a single photon-like resonance, a phenomenon sometimes described as the high-power cavity magnon-polaritons regime [60]. For more theoretical analysis and validation under both linear and nonlinear regimes, refer to the Supplementary Materials. These results collectively demonstrate the model’s capability in predicting the intricate hybrid quantum coupling, especially in complex, realistic integrated chip setup.

### v2

@@ Abstract
Simulating hybrid magnonic quantum systems remains a challenge due to the large disparity between the timescales of the two systems. We present a massively parallel GPU-based simulation framework that enables fully coupled, large-scale modeling of on-chip magnon-photon circuits. Our approach resolves the dynamic interaction between ferromagnetic and electromagnetic fields with high spatiotemporal fidelity. To accelerate design workflows, we develop a physics-informed machine learning surrogate trained on the simulation data, reducing computational cost while maintaining accuracy. This combined approach reveals real-time energy exchange dynamics and reproduces key phenomena such as anti-crossing behavior and the suppression of ferromagnetic resonance under strong electromagnetic fields. By addressing the multiscale and multiphysics challenges in magnon–photon modeling, our framework enables scalable simulation and rapid prototyping of next-generation quantum and spintronic devices.
@@ 1 Introduction
Hybrid quantum systems, which combine distinct physical platforms, are a promising route toward advanced quantum technologies, as they harness strong interactions that may not be readily achievable in a single platform [Nielsen_Chuang_2010, 2015Gershon]. These systems take many forms, coupling any two (or more) quantum platforms — for example, superconducting qubits  [Scarlino2019, Tabuchi2015], microwave resonators  [Clerk2020], single spins [Viennot2015], spin ensembles  [Tabuchi2015, Hou2019, Li2019, Chen2018], or mechanical resonators [o2010phonon, andrews2015phonon, manenti2017phonon] — to harness strong interactions. These heterogeneous systems leverage complementary advantages of each component, but their rich multi-physics interactions pose formidable modeling challenges. A prominent example is cavity magnonics, where collective spin excitations (magnons) couple with microwave photons in a resonant cavity to form hybrid magnon–polariton modes when tuned into resonance [Zhang2014, ZareRameshti2022, Awschalom2021]. While an avoided crossing indicates hybridization, it is not sufficient for strong coupling; the system enters the strong-coupling regime only when the coupling rate $g$ (half the on‑resonance splitting) exceeds the dissipation rates of the uncoupled magnon and photon modes, in which case energy oscillates coherently between them and the hybrid modes exhibit remarkable coherence owing to the high $Q$ factors of microwave resonators and the narrow ferromagnetic resonance (FMR) linewidth [Match2019]. These states are essential for quantum operations such as mode swapping [Match2019, zhuang2024dynamical], quantum state storage  [Zhang2015, Tabuchi2015, Pishehvar2024], and dynamic control of energy exchange  [Xu2020, Pishehvar2024]. The hallmark experimental signature of strong magnon–photon coupling is a pronounced avoided crossing (mode splitting) in the frequency spectrum, in agreement with theoretical predictions [MagnonPhotonTheory] and observed in many 3D  [Zhang2014, Goryachev2014] and on-chip 2D [Huebl2013, Hou2019, Li2019] cavity based systems. In the time domain, the coupled system exhibits Rabi-like oscillations with very high energy exchange per cycle, reflecting the complete hybridization of the two excitations into a single quantum system [zhuang2024dynamical]. Achieving practical quantum operations with magnon-photon systems requires precise dynamical control over energy transfer between the two hybrid magnon-polariton modes [Boventer2020, Golovchanskiy2021]. Therefore, developing quantum devices that leverage these effects necessitates a detailed understanding of their time-domain dynamics.
@@
However, accurately modeling the complex dynamics of such hybrid systems presents a significant computational challenge. Existing modeling approaches have been limited in scope. Most analytical treatments require inputting an assumed magnon–photon coupling strength, and the few numerical studies to date have often resorted to reduced dimensions (1D or 2D models) or macrospin approximations to mitigate computational cost [MagnonPhotonTheory]. Conventional micromagnetic modeling strategies rely on the magnetostatic approximation [donahue1999oommf, mumax3, bjork2021magtense], which has provided valuable insights, but which typically neglects the dynamic nature of electromagnetic (EM) fields, including time-dependent displacement currents. The underlying physics of magnon-photon coupling spans multiple domains and scales, requiring a multiphysics simulation that solves Maxwell’s equations for electromagnetic fields alongside the Landau–Lifshitz–Gilbert (LLG) equation for magnetization dynamics [zhuang2024dynamical]. Solving these coupled equations self-consistently in three dimensions with realistic material parameters is computationally intensive, often necessitating high-performance computing (HPC) resources to achieve adequate spatial and temporal resolution [Yao2022]. The disparate time scales and length scales involved – from microwave photons to nanometer-scale spin precessions – further complicate direct simulation. Maxwell-LLG coupling is achieved by self-consistently updating the dynamic EM fields in sync with magnetization dynamics, thereby capturing the bidirectional interaction between microwave photons and magnons  [Yao2022, Yao2021, Yao2018, zhuang2024dynamical, xu2024slow, pishehvar2025demand, maksymov2015rigorous, yu2020circulating]. The majority of such time-domain dynamical simulations employ either explicit schemes like finite-difference time-domain (FDTD) leapfrog integrators  [Yao2022, maksymov2015rigorous, zhuang2024dynamical, pishehvar2025demand, xu2024slow], or implicit schemes such as alternating-direction implicit (ADI) FDTD and Crank–Nicolson methods  [Yao2021, Yao2018]. While explicit FDTD is straightforward, it is only conditionally stable (restricted by the Courant limit), leading to significant oversampling of the EM wave in the time domain. To ease numerical constraints, researchers have explored strategies such as artificially increasing the medium permittivity—an approach that relaxes time-step oversampling [zhuang2024dynamical]. However, this strategy might be difficult to extend to complex device structures with highly inhomogeneous local permittivity and conductivities, as in on-chip devices. On the other hand, implicit methods offer improved (often unconditional) stability, but they require solving more complex update steps. Additionally, implementing precise boundary conditions in implicit mathematical frameworks, such as perfectly matched layers (PML), is challenging and limits scalability in real-world quantum devices  [Yao2018]. HPC techniques, including GPU acceleration and parallel FDTD-LLG solvers combined with domain decomposition, significantly enhance simulation speed and scalability, enabling large-scale modeling on modern supercomputing architectures  [fu2015finite, Yao2022]. Finite-element time-domain models that rely on commercial software have also been adopted, but they suffer from the overhead of assembling and solving large linear systems at each time step, leading to high CPU and memory usage [yu2020circulating]. Nonetheless, to the best of the authors’ knowledge, dynamical time-domain simulation of practical sized, on-chip hybrid magnon–photon circuits has remained elusive [Hou2019, Li2019]. This underscores the need for more powerful computational methods to simulate hybrid magnon–photon systems in full detail.
@@
The main numerical challenges arise from the need for large-scale three-dimensional (3D) simulations and fine meshes to accurately capture the detailed structures. In many practical scenarios, however, only a few spatial observation points are needed for circuit analysis tasks. This discrepancy underscores the potential for more efficient modeling strategies that focus resources on critical regions and time windows, reducing overall costs while preserving key physics. Machine learning (ML) has emerged as a powerful approach to tackle the computational challenges posed by high-dimensional and nonlinear dynamical systems [brunton2020machine, wang2023scientific]. In particular, frameworks such as Physics-Informed Neural Networks (PINNs) [karniadakis2021physics, raissi2019physics], Neural Operators (e.g., DeepONet  [lu2021learning] and Fourier Neural Operator [li2020fourier]), and Neural Discrete Equilibrium (NeurDE) [neurde_TR] have demonstrated strong capabilities in modeling partial differential equations (PDEs), by efficiently capturing their underlying spatial-temporal dynamics. Furthermore, recent advances in sequence modeling architectures, including Recurrent Neural Networks (RNNs) [hochreiter1997long], state-space models (SSMs) [gu2022s4_iclr, annan_tuning_TR] and Transformers [vaswani2017attention], have shown remarkable success in predicting complex time-evolving physical phenomena [Tang2022, Ravuri2021, rao2023encoding, bi2023accurate, lam2023learning, song2024forecasting, lyu2024wavecastnet, learnConservation1_TR].
@@ 2 Results
Figure 1: Magnon-photon hybrid quantum circuit modeling setup and simulated field profiles. (a) Perspective view of the magnon-photon coupling circuit, comprising a CPW resonator and a ferromagnetic thin film. The CPW resonator consists of a center conductor and lateral ground conductors (yellow) positioned on a silicon substrate (blue). The ferromagnet, shown as a red stripe, is placed on top of the center conductor. Material and geometrical parameters are detailed in the main text. (b)–(d) Cross-sectional views of the circuit, overlaid with the simulated magnetic field (H) distribution at different locations along the longitudinal direction of the CPW resonator, as indicated in (a). (e)–(g) Top views of the circuit, overlaid with the simulated electric and magnetic field distributions in the z-plane at different observation points, as marked in (a). (h) Theoretical field distributions along the longitudinal direction of the circuit, where the electric field $E_{x}$ is concentrated in the lateral air gap, and the magnetic field $H_{x}$ is observed above and beneath the center conductor.
@@
We employ a fully coupled dynamic Maxwell–LLG solver, ARTEMIS, based on the approach in [Yao2022], using an explicit FDTD leap-frog scheme to update the electric field E, magnetic field H, and magnetization M. In nonmagnetic regions, only Maxwell’s equations are solved (with M  set to zero). In magnetic regions, Maxwell’s equations and the LLG equation are solved self-consistently in a mutual feedback loop (see Numerical Model. in Methods for details). All field variables computed by the solver—E, H, and M—achieve second-order accuracy, owing to the trapezoidal time-domain discretization of the LLG equation and the leap-frog time-marching scheme for Maxwell’s equations. To handle large problem sizes, we implement both multicore and GPU parallelization for the spatial mesh, as illustrated in Figure 2(a). We use the AMReX [AMReX_JOSS] library to partition the computational domain into non-overlapping grids, each assigned to an MPI rank. AMReX follows an MPI+X parallelization model, where X can be OpenMP (for multicore CPUs) or CUDA (for GPUs).
@@
The input data for the physics-informed ML model is collected from nine discrete spatial locations, shown in Figure 2(b). These points lie in the middle of the ferromagnet layer along its thickness and include the film’s center, as well as the four corners and midpoints of the four edges in the lateral plane. More details can be found in the ML Training Implementation. section of Methods. Figure 2(f) illustrates the general ML framework for sequential modeling of the magnetization. The model is based on the Long Expressive Memory (LEM) module [rusch2022iclr], and it comprises an LEM encoder and decoder. The structure of the LEM cell is illustrated in Figure 2(d), which shows the computation flow across its internal states. At each time step, the input $X_{t}$, the previous hidden state $H_{t-1}$ and latent state $C_{t-1}$ are linearly projected and combined to generate two adaptive time-step parameters: $\Delta t$ for updating the latent state $C_{t}$, and $\overline{\Delta t}$ for updating the hidden state $H_{t}$. These two time-step parameters act as gating coefficients, controlling the extent to which newly computed nonlinear activations (via $\tanh$) are blended into each state. The latent state $C_{t}$ is updated first, capturing fast-evolving features from the input and hidden state. This updated $C_{t}$ is then transformed and used to update the hidden state $H_{t}$, which serves as the main memory for downstream prediction. Additional implementation details and the mathematical formulation are provided in the LEM Cell. of Methods. The LEM module processes the input sequence (200 temporal samples) using a stacked encoder composed of two LEM cells, which evolve coupled hidden $H$ and latent states $C$ with adaptive time constants. The final states from the encoder are used to initialize a two-stage decoder, which begins from a zero input and autoregressively generates a fixed-length output sequence. At each decoding step, the internal states are updated using the same LEM dynamics, and the top-layer hidden state is passed through a fully connected layer to produce the output. The model is trained using a composite loss function that integrates three component:
@@
Given the complexity of the hybrid quantum system, we incorporated Curriculum Learning [2101.10382] to improve generalization and help the model better capture long-term dynamics and the underlying physics. As shown in Figure 2(e), this approach enables the model to handle tasks of increasing complexity over time. During training, three key hyperparameters are dynamically adjusted:
@@
Training begins with simpler scenarios—shorter prediction sequences, no physics loss, and a higher learning rate—and gradually moves to more complex cases involving longer sequences, physics loss, and a lower learning rate. This staged approach ensures that the model generalizes effectively while capturing the intricate physics governing the field interactions. Details on how the training and testing datasets were curated are provided in the Datasets Curation. section of Methods.
Figure 2: The HPC and ML frameworks for magnon-photon hybrid quantum modeling. (a) Simulation domain setup and GPU partitioning in the Maxwell–LLG coupled numerical model, ARTEMIS. The GPU parallelization leverages the grid-structured GPU library AMReX [AMReX_JOSS], achieving nearly flat weak-scaling performance. (b) Top view of the magnon-photon hybrid quantum circuit; the cross marks indicate probing points used for data collection. (c) Example time-series data of the magnetization M recorded at a single probing point. (d) Detailed architecture of the LEM cell. The cell updates a latent state ($C_{t}$) and a hidden state ($H_{t}$) using the input $X_{t}$, previous states, and learnable time constants ($\Delta t$). $C_{t}$ captures short-term dynamics, while $H_{t}$ integrates transformed latent features to maintain temporal context. (e) The curriculum learning strategy used to train the ML model, with a staged schedule with progressively adjusted learning rates, batch sizes, sequence lengths, and physics-based loss weights to stabilize training and enhance convergence. (f) Overall ML framework for time-series prediction, built out of the LEM-based architecture. The encoder consists of sequential LEM Cells that process the input sequence (X) into two evolving internal states: the latent state (C); and the hidden state (H). These states are used to initialize an auto-regressive decoder, which also consists of stacked LEM Cells. At each step, the decoder outputs a hidden state (H) that is passed through a shared fully connected (FC) layer to produce the predicted output.
@@
Our Maxwell-LLG dynamical modeling has predicted the dynamically evolving and interacting electric field and magnetization, shown in Figure 3. Figure 3(a) presents top-view snapshots of $E_{x}$ (top row) and $M_{x}$ (bottom row) at selected time instants, with an applied magnetic DC bias $H_{0}=2050$ Oersted. According to the Kittel’s law, under such $H_{0}$, the FMR frequency is $f_{\textrm{m}}=\omega_{\textrm{m}}/2\pi=(\gamma/2\pi)\sqrt{H_{0}(H_{0}+4\pi M_{S})}=15.027$ GHz, detuned from the photon frequency $f_{\textrm{p}}$ by about $\Delta=200$ MHz. This small detune enables strong mode hybridization and, consequently, strong magnon-photon coupling. The $E_{x}$ fields are observed at the same location described in Figure 1(e), i.e., along the vertical center plane passing through the CPW resonator’s lateral air gap. The $M_{x}$ fields are observed at the same location described in Figure 1(f), i.e., on the top surface of the ferromagnetic stripe. For additional details on these observation points, see Figure 1(a). At 0.0235 ns, the electric field appears only in the feeding patches, without exciting resonant modes (photons) in the CPW resonator, and no dynamic magnetization is observed in the ferromagnet. By 0.3528 ns, both $E_{x}$ and $M_{x}$ emerge in the resonator. Starting at 0.9819 ns, even though the input field excitation has vanished (the input signal completely decays by 0.5 ns, as shown in Figure 3(b)), oscillations of $E_{x}$ and $M_{x}$ persist within the CPW resonator. Notably, $E_{x}$ peaks near the resonator’s ends, whereas $M_{x}$ peaks at the center. Both fields exhibit well-defined oscillations with only minor decay, indicating low damping—a result of minimal Gilbert damping, small metallic losses, and limited electromagnetic leakage into air or substrate. Furthermore, $E_{x}$ and $M_{x}$ alternate in reaching their standing-wave maxima: when $E_{x}$ is at a peak, $M_{x}$ is at a node, and vice versa. This phase alternation underscores strong energy exchange between the magnon and photon polaritons, more clearly seen in the transient response of Figure 3(c), and the yielded mode splitting in the Fourier transformed $E_{x}$ spectra shown in Figure 3(d). Figure 3(e) visualizes the spectra of $E_{x}$ magnitude under various $H_{0}$, revealing a clear anti-crossing behavior, indicating coherent coupling between magnons and photons [ZareRameshti2022]. The simulated spectrum matches theoretical analysis as shown in Extended Data Figure 1. The $S$-parameter ($|S_{12}|$) spectra shown in Figure 3(f) is converted from the $E_{x}$ field following the workflow based on field calculation in  [Sawant2023], indicating the transmission rate from one end to the other of the microwave resonator, also showing an anti-crossing spectra. Both figures show mode splitting $g_{\textrm{mp}}=670$ MHz, where $g_{\textrm{mp}}$ is the total magnon-photon coupling strength, defined as $g_{\textrm{mp}}=g_{0}\sqrt{N}$, with $g_{0}$ defined as the coupling strength between photons and individual spins in the ferromagnet, and $N$ is the number of spins in the ferromagnet [MagnonPhotonTheory, Zhang2014]. $g_{0}$ contributes to the total Hamiltonian of the system:
@@
One unique advantage of numerical simulation is the ability to generate signals on demand, including those with intensities that are experimentally unattainable. Extended Data Figure 2 displays the electric-field spectra as the excitation intensity is increased. It demonstrates that, in the absence of significant spin-spin interactions (e.g., exchange coupling), progressively stronger microwave magnetic fields suppress the magnon dynamics and leave only the cavity photon mode. Such behaviour arises because, at high powers, nonlinear spin-wave interactions—particularly three-magnon splitting—induce a Suhl instability that transfers energy from the uniform ferromagnetic resonance to non-uniform spin waves; this saturation of the Kittel mode causes the magnon–photon anticrossing gap to close and yields a single photon-like resonance, a phenomenon sometimes described as the high-power cavity magnon-polaritons regime [nonlinear2023]. For more theoretical analysis and validation under both linear and nonlinear regimes, refer to the Supplementary Materials. These results collectively demonstrate the model’s capability in predicting the intricate hybrid quantum coupling, especially in complex, realistic integrated chip setup.